RMC issued perpetual preferred stock with a 20% annual dividend. The stock currently yields 4%, and its par value is $100. Suppose interest rates rise and pull the preferred stock’s yield up to 10%. How much would its market value change?
1
Expert's answer
2020-05-11T17:56:00-0400

100*0.2=20-annual dividend

20/0.04=500-currently market value

20/0.1=200-new market value

market value of stock change=200-500=-300$